# LLM Collaboration With Multi-Agent Reinforcement Learning

### v1

@@ Abstract
A large amount of work has been done in Multi-Agent Systems (MAS) for modeling and solving problems with multiple interacting agents. However, most LLMs are pretrained independently and not specifically optimized for coordination. Existing LLM fine-tuning frameworks rely on individual rewards, which require complex reward designs for each agent to encourage collaboration. To address these challenges, we model LLM collaboration as a cooperative Multi-Agent Reinforcement Learning (MARL) problem. We develop a multi-agent, multi-turn algorithm, Multi-Agent Group Relative Policy Optimization (MAGRPO), to solve it, building on current RL approaches for LLMs as well as MARL techniques. Our experiments on LLM writing and coding collaboration demonstrate that fine-tuning MAS with MAGRPO enables agents to generate high-quality responses efficiently through effective cooperation. Our approach opens the door to using other MARL methods for LLMs and highlights the associated challenges.
Introduction
@@
Leveraging billions of parameters and extensive pre-training on large-scale datasets, state-of-the-art LLMs have demonstrated remarkable capabilities across diverse domains (Grattafiori et al. 2024; Achiam et al. 2023; Anil et al. 2025). To adapt to specific applications or align with human preferences, fine-tuning has emerged as a critical secondary training stage. Compared to supervised fine-tuning, Reinforcement Learning (RL) enables more generalizable learning for complex, multi-turn tasks through human-aligned reward design, making it an important technique for fine-tuning (Ouyang et al. 2022; Guo et al. 2025; Ziegler et al. 2020).
@@
Likewise, Multi-Agent Systems (MAS) have been extensively studied over the past decades, with substantial progress in modeling and solving problems involving multiple agents (Littman 1994; Shoham and Leyton-Brown 2009; Stone and Veloso 2000). In particular, advances in cooperative MAS have demonstrated strong potential for enabling effective collaboration in distributed settings, such as games, robotics, and traffic control (Samvelyan et al. 2019; Vinyals et al. 2017; Berner et al. 2019; Amato et al. 2016; Wiering 2000). These developments motivate the application of MAS principles and techniques to LLM collaboration, where multiple LLMs working together can solve more complex tasks in a more robust and efficient manner.
@@
There has been some recent work on coordinating multiple LLMs. Some approaches implement coordination at the inference stage, enabling agents to interact through debate, discussion, or verification (Du et al. 2023; Wu et al. 2023a; Lifshitz, McIlraith, and Du 2025). These methods operate at the prompt level, with fixed models that are not tuned toward coordination-centric objectives. The agents may have conflicting answers or spread incorrect information to other participants, limiting performance (Cemri et al. 2025; Estornell and Liu 2024). Moreover, the design of effective prompts remains difficult and unclear. Other approaches fine-tune agents independently with individual or role-conditioned rewards. However, they require carefully curated rewards for each individual or role (Slumbers et al. 2024; Liu et al. 2025; Subramaniam et al. 2025), and, as independent learning methods, lack convergence guarantees (Tan 1993).
@@
In this paper, we model LLM collaboration as a cooperative MARL problem (Albrecht, Christianos, and Schäfer 2024) and formalize it as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) (Oliehoek and Amato 2016). In LLM collaboration, multiple trainable LLMs generate responses synchronously based on their individual prompts. The external environment evolves according to the joint responses until the dialog ends. This general model allows a wide range of problems to be modeled and solved using versions of MARL algorithms. Following the efficient practice of Group Relative Policy Optimization (GRPO) (Guo et al. 2025), we propose Multi-Agent GRPO (MAGRPO) that trains LLMs in a multi-turn setting. MAGRPO leverages centralized group-relative advantages for joint optimization, while preserving decentralized execution for each agent. The resulting method builds off of state-of-the-art LLM approaches in GRPO and MARL approaches for centralized training and decentralized execution, such as MAPPO (Yu et al. 2022). Our experiments demonstrate that MAGRPO develops various LLM cooperation schemes, improving response efficiency with high quality.
@@
Our contributions can be summarized as follows: (i) We model the LLM collaboration as a cooperative MARL problem, where multiple LLMs cooperate to generate joint responses; (ii) We implement the MAGRPO algorithm, which optimizes agent cooperation through aligned rewards while maintaining decentralized execution to maintain efficiency; (iii) Our experiments demonstrate that fine-tuning with MAGRPO improves both response efficiency and quality in writing and coding collaboration; (iv) We provide a detailed analysis of the limitations of existing approaches and outline open challenges in applying MARL to LLM collaboration.
Related Work
Test-Time Multi-Agent Interaction
@@
Recent work employs multiple agents with specialized roles interacting through diverse pipelines at test-time to enhance response quality. In multi-agent debate, agents iteratively formulate positions by reviewing other agents’ outputs, where the final decision or answer is determined by majority voting or a summarizer (Du et al. 2023; Chan et al. 2023; Liang et al. 2024). Role-based approaches allocate tasks across specialized agents (Wu et al. 2023a; Qian et al. 2024; Hong et al. 2024). For example, an agent may function as a verifier to assess the correctness of outputs (Skreta et al. 2023; Lifshitz, McIlraith, and Du 2025; Setlur et al. 2025), while another may act as a macro-planner to orchestrate workers’ responses. However, these multi-agent frameworks rely on prompt-level interactions among agents, often leading to ineffective communication and computational inefficiency. Moreover, the design of effective prompts and role assignment remains unclear, as prompts usually fail to reliably guide agent behavior, enforce role adherence, or support coherent coordination across tasks. These limitations motivate us to fine-tune LLMs in MAS to improve their cooperation.
Multi-Agent Fine-Tuning
@@
Recent work has explored fine-tuning LLMs to improve their performance across diverse domains, e.g., arithmetic reasoning, navigation, and hidden-role games (Ma et al. 2025; Slumbers et al. 2024; Sarkar et al. 2025). These approaches typically employ individual rewards or rewards conditioned on specific roles (Park et al. 2025; Liu et al. 2025; Subramaniam et al. 2025). Such reward structures often require careful manual specification, and their underlying rationale is rarely well justified. The misaligned or conflicting incentives can hinder effective coordination. Moreover, these methods lack convergence guarantees, as each agent learns independently in a non-stationary environment where other agents are simultaneously updating their policies. In this paper, we focus on cooperative scenarios, where LLMs are jointly trained with interpretable, human-aligned rewards.
Cooperative MARL for LLM Collaboration
@@
As discussed by a number of companies (NVIDIA 2024; Anthropic 2024), a team of agents could be used to generate a complex codebase. The code would be difficult, costly, and time-consuming to generate with a single agent, but a group of LLMs could do so quickly and cheaply. None of these agents is self-interested, but they are trainable in a scheme such as the one discussed below. Using a joint reward allows agents to specialize as needed to complete the task without complex prompt or reward engineering.
Problem Formalization
We formalize collaboration among LLMs as a subclass of the cooperative MARL problem, considering LLM agents and the types of problems they are solving. This problem is a form of a Dec-POMDP (Oliehoek and Amato 2016), which allows cooperation through a joint reward while preserving scalable decentralized control. We show 2 instantiations of our framework in writing and coding tasks in the experiments section.
@@
In Dec-POMDP, since the states are not directly observed, each agent maintains its local observation-action history $\mathbf{h}=\{h_{1},\cdots,h_{n}\}$ to infer information about state. A solution to a Dec-POMDP is a joint policy that maximizes the expected cumulative reward, $\boldsymbol{\pi}^{*}=\{\pi_{1}^{*},\cdots,\pi_{n}^{*}\}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{H-1}R(s^{\text{acc}}_{t},\mathbf{a}_{t})\right]$. A joint polict is a set of local policies $\pi_{i}$, which conditions on the local observation-action history $h_{i,t}=\{o_{i,0},a_{i,0},\ldots,o_{i,t}\}$.
RL methods for Dec-POMDPs have become a popular topic (e.g., (Foerster et al. 2024; Lowe et al. 2020; Foerster et al. 2018; Rashid et al. 2018; Wang et al. 2021; Yu et al. 2022; Albrecht, Christianos, and Schäfer 2024)) with methods successful at scaling to large state, action and observation spaces. Many methods use Centralized Training for Decentralized Execution (CTDE), where they use some centralized information during training (e.g., a centralized value function estimate) but are still able to execute in a decentralized manner when training is complete.
Challenges in LLM Collaboration
@@
Representations in Natural Language
Unlike traditional cooperative MARL agents, LLM agents operate over natural language, receiving instructions and generating responses as sequences of tokens. MARL approaches could model this problem at the token or prompt/response level. At the token level, the number of actions and observations is smaller, but the problem horizon can be very long. At the prompt/response level, the actions and observations space is much larger, but the horizon is much shorter. Moreover, token-level rewards are often uninformative, as both queries and responses must form coherent and semantically meaningful structures. As adopted in prior RL methods (Ouyang et al. 2022; Guo et al. 2025; Rafailov et al. 2024), we model each LLM agent’s decision-making process as a direct mapping from input instructions to complete responses to enable efficient and stable training. Nevertheless, the best modeling and solution approaches remain an open question.
Training Paradigm
As mentioned above, many MARL methods use centralized training for decentralized execution (CTDE). Unfortunately, standard CTDE methods use centralized value models in the form of centralized critics (Foerster et al. 2024; Lowe et al. 2020; Yu et al. 2022) or mixers in value decomposition methods (Rashid et al. 2018; Wang et al. 2021). Such architectures allow additional information and coordination during training but do not scale well to very large action and observation spaces (such as those in our problem). Conversely, Decentralized Training and Execution (DTE) methods (Amato 2025) train a set of models, one for each agent in a decentralized manner. DTE approaches are typically more scalable but don’t use additional information during training (even when it is available). It is an open question which paradigm to use to maximize performance while maintaining scalability in the LLM collaboration problem. In this paper, we balance decentralized execution with centralized training using group-based Monte Carlo estimates. Experiments show the effectiveness of our approach on short-horizon tasks.
MAGRPO
@@
12:   Estimate $\widehat{A}^{(g)}_{t}$, $\forall g\in\mathcal{G}$ according to Equation 1
13:   Calculate $J(\theta_{i})$, $\forall i\in\mathcal{I}$ according to Equation 2
@@
Algorithm 1 shows the procedure of MAGRPO. Given a dataset $\mathcal{D}$ containing task information (e.g., the descriptions of coding problems), $n$ LLMs are optimized, each with a policy parameterized by $\theta_{i}$ and guided by a reward model $R$. In each episode, a task is sampled from the given dataset $\mathcal{D}$, which is used to construct initial observations $\mathbf{o}_{0}=\{o_{1,0},\cdots,o_{n,0}\}$ and histories $\mathbf{h}_{0}=\{h_{1,0},\cdots,h_{n,0}\}$. Taking inspiration from the single-agent GRPO algorithm (Guo et al. 2025), at each turn $t$, each agent takes action by generating a group of responses $a^{\mathcal{G}}_{i,t}=\{a^{(1)}_{i,t},\cdots,a^{(G)}_{i,t}\}$ following its policy $\pi_{i}(\cdot|h_{i,t}^{\mathcal{G}})$ based on its observation-action history $h_{i,t}^{\mathcal{G}}=\{h^{(1)}_{i,t},\cdots,h^{(G)}_{i,t}\}$. The actions of individual agents are aggregated to form a group of joint actions $\mathbf{a}^{\mathcal{G}}_{t}=\{a^{\mathcal{G}}_{0,t},\cdots,a^{\mathcal{G}}_{n,t}\}$. The agents receive a group of joint rewards $r^{\mathcal{G}}_{t}$ for their responses $\mathbf{a}^{\mathcal{G}}_{t})$, which also conditions on the accessible part of the state $R(\cdot|s^{\text{acc},\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$. The joint actions triggers the transition $T(\cdot|s^{\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$, where agents receive new observations $o^{\mathcal{G}}_{i,t+1}=\{o^{(1)}_{i,t},\cdots,o^{(G)}_{i,t}\}$ and use them to construct histories $h^{\mathcal{G}}_{i,t+1}=\{h^{\mathcal{G}}_{i,t},a^{\mathcal{G}}_{i,t},o^{\mathcal{G}}_{i,t+1}\}$. This process continues until terminated at turn $H$.
@@
We employ stochastic gradient descent to train agents at the end of each episode. Without explicit value models, estimating history-action values from a single rollout incurs high variance. To stabilize training, we estimate the expected return of the current state by averaging over a group of Monte Carlo samples $\{R^{(1)}_{t},\cdots,R^{G}_{t}\}$. As a result, we are able to generate a centralized estimate (which is common in MARL) without a large value model. For each turn $t$, the advantage of each joint action in the group is calculated as,
$\widehat{A}^{(g)}_{t}=\frac{R^{(g)}_{t}-\frac{1}{G}\sum_{g=1}^{G}R^{(g)}_{t}}{\sigma(R^{(g)}_{t})}$, (1)
where $\sigma(R^{\mathcal{G}}_{t})$ represents the standard deviation of a group of expected returns, and $R^{(g)}_{t}=\sum_{\tau=t}^{H-1}r^{(g)}_{\tau}$.
@@
Inspired by GRPO (Guo et al. 2025) and MAPPO (Yu et al. 2022), the centralized advantage values can be used to update policy $\pi_{i}$ (parameterized by $\theta_{i}$) for each agent $i$,
@@
$J($ $\theta_{i})=\mathbb{E}_{\mathbf{o}_{0}\sim\mathcal{D},\mathbf{h}^{\mathcal{G}}\sim\boldsymbol{\pi}_{\boldsymbol{\theta},\textbf{old}}}$ (2)
$\Bigg{[}\frac{1}{|B|}\frac{1}{|\mathcal{G}|}\sum_{h_{i}^{\mathcal{G}}\in B}\sum_{g\in\mathcal{G}}\min\left(\rho^{(g)}_{i,t}\widehat{A}^{(g)}_{t},\text{$\varepsilon$-clip}(\rho^{(g)}_{i,t})\widehat{A}^{(g)}_{t}\right)\Bigg{]}$,
where $\rho^{(g)}_{i,t}=\frac{\pi_{\theta_{i}}(a^{(g)}_{i,t}|h^{(g)}_{i,t})}{\pi_{\theta_{i},\text{old}}(a^{(g)}_{i,t}|h^{(g)}_{i,t})}$ denotes the importance sampling ratio between the updated and previous policies.
Experiments
We evaluate MAGRPO on LLM writing and coding collaboration. Datasets, reward specifications, and additional results are provided in the Appendix.
Writing Collaboration
Method Dataset Efficiency Article Quality (%) Return (%)
Speed Response Time Structure Consistency Coherence
Single Model TLDR 64.1 6.6 43.8 97.6 52.8 36.7
arXiv 65.4 6.5 51.2 87.2 71.1 44.9
Parallel Generation TLDR 185.6 2.1 25.9 98.3 56.5 23.2
arXiv 190.6 2.1 71.5 64.2 61.5 59.6
Sequential Generation TLDR 98.7 4.3 33.5 98.5 64.5 21.7
arXiv 85.8 4.3 92.4 97.8 64.3 87.7
One-Round Discussion TLDR 100.4 4.3 35.9 98.8 60.8 22.3
arXiv 95.4 4.3 84.6 71.8 66.0 76.6
MAGRPO (Ours) TLDR 202.3 2.1 98.7 97.1 78.5 94.5
arXiv 193.8 2.1 97.9 96.2 69.7 93.1
Table 1: Performance of MAGRPO against baselines on TLDR and arXiv. Speed (tokens/s) and response time (s) are measured on GeForce RTX 5090s. Results are normalized within the return scale. Bolds indicate the best performance on each dataset.
We explore LLM collaboration for article writing using MAGRPO across 2 classic tasks: summarization and expansion.
(a) TLDR summarization
(b) arXiv abstract expansion
Figure 2: Normalized returns on writing collaboration: (a) structural wellness (dashed green); (b) style consistency (dashed red); (c) coherence (dashed orange); (d) total rewards (solid blue). All returns are normalized within the return scale.
TLDR Summarization
When reading a long article, readers often seek to quickly grasp its core idea. If the topic is of interest, they may wish to delve deeper into specific details while still avoiding a complete reading through the full document. This calls for a summarization system to generate summaries at varying levels of detail. We frame this task using TLDR summarization as an illustrative example.
The TLDR dataset comprises unabridged Reddit posts in the prompt and concise summaries appended by the author in the completion. In our experiment, 2 Qwen3-1.7B agents independently summarize the prompt without using completion. The first agent functions as a core-idea (TLDR) generator, producing a concise paragraph, while the second agent serves as a detailed summarizer, providing more comprehensive information.
To quantify the summarization quality, we employ a relatively simple combination of 3 metrics. Structure measures the lengths and the length ratio of the two summaries, to ensure the TLDR is concise and the detailed summary is sufficiently long. Style consistency is assessed using the normalized Jaccard similarity coefficient, calculated as the ratio of the intersection size to the union size of unique words (or n-grams) between responses. A high style consistency reward typically indicates that the summarizers adopt similar stylistic patterns while avoiding identical wording. Logical coherence is quantified by counting the occurrences of transition words. Positive reward is given for using transition words, but the reward decreases logarithmically as more are used. These metrics are simple approximations of what more complex reward models may evaluate. Other (simpler or more complex) metrics or reward models could also be used. The total reward combines these metrics through a weighted summation. More details regarding our reward model and hyperparameters are provided in the Appendix.
arXiv Expansion
Writing a long article typically requires contributions from multiple writers, each responsible for different sections. As a simple scenario, 2 agents can collaborate to generate introductions from the abstract of arXiv papers. The first agent outlines the research background and motivation, while the other presents the proposed methods and their experiments. The combined paragraphs should be coherent and consistent in style. Similar to the reward model in TLDR summarization, we employ the same evaluation metrics as proxies, with threshold hyperparameters specifically adjusted for this task.
Baselines
We adopt a single-agent model and 3 representative multi-agent methods as our baselines. To minimize the influence of prompts on our comparison, we keep the task description fixed and only add minimal coordination instructions. Specifically, for the single-agent baseline, we prompt with the article to be manipulated, the agent’s role (summarizer or expanding writer), and specific user instructions (e.g., format requirements). Naive concatenation builds on it by dividing the task into subtasks, assigning each agent a specific portion to complete in parallel without explicit communication. The sequential pipeline introduces one-way communication, allowing one agent to respond based on both the task description and the other agent’s output. The one-round discussion baseline enables bidirectional communication: agents first receive the same prompts as in naive concatenation, then the prompts are augmented with the other’s first-turn response in the second turn. All baseline methods operate without fine-tuning and depend solely on prompt-level interactions. Detailed prompts for each baseline are provided in the Appendix.
(a) Single-Turn MAGRPO on HE
(b) Multi-Turn MAGRPO on HE
(c) Single-Turn MAGRPO on CHE
(d) Multi-Turn MAGRPO on CHE

### v2

@@ Abstract
A large amount of work has been done in Multi-Agent Systems (MAS) for modeling and solving problems with multiple interacting agents. However, most LLMs are pretrained independently and not specifically optimized for coordination. Existing LLM fine-tuning frameworks rely on individual rewards, which require complex reward designs for each agent to encourage collaboration. To address these challenges, we model LLM collaboration as a cooperative Multi-Agent Reinforcement Learning (MARL) problem. We develop a multi-agent, multi-turn algorithm, Multi-Agent Group Relative Policy Optimization (MAGRPO), to solve it, building on current RL approaches for LLMs as well as MARL techniques. Our experiments on coding collaboration demonstrate that fine-tuning MAS with MAGRPO enables agents to generate high-quality responses efficiently through effective cooperation. Our approach opens the door to using other MARL methods for LLMs and highlights the associated challenges.
1 Introduction
@@
Leveraging billions of parameters and extensive pre-training on large-scale datasets, state-of-the-art LLMs have demonstrated remarkable capabilities across diverse domains Grattafiori et al. (2024); Achiam et al. (2023); Anil et al. (2025). To adapt to specific applications or align with human preferences, fine-tuning has emerged as a critical secondary training stage. Compared to supervised fine-tuning, Reinforcement Learning (RL) enables more generalizable learning for complex, multi-turn tasks through human-aligned reward design, making it an important technique for fine-tuning Ouyang et al. (2022); Guo et al. (2025); Ziegler et al. (2020).
@@
Likewise, Multi-Agent Systems (MAS) have been extensively studied over the past decades, with substantial progress in modeling and solving problems involving multiple agents Littman (1994); Shoham and Leyton-Brown (2009); Stone and Veloso (2000). In particular, advances in cooperative MAS have demonstrated strong potential for enabling effective collaboration in distributed settings, such as games, robotics, and traffic control Samvelyan et al. (2019); Vinyals et al. (2017); Berner et al. (2019); Amato et al. (2016); Wiering (2000); Liu et al. (2022). These developments motivate the application of MAS principles and techniques to LLM collaboration, where multiple LLMs working together can solve more complex tasks in a more robust and efficient manner.
@@
There has been some recent work on coordinating multiple LLMs. Some approaches implement coordination at the inference stage, enabling agents to interact through debate, discussion, or verification Du et al. (2023); Wu et al. (2023a); Lifshitz et al. (2025). These methods operate at the prompt level, with fixed models that are not tuned toward coordination-centric objectives. The agents may have conflicting answers or spread incorrect information to other participants, limiting performance Cemri et al. (2025); Estornell and Liu (2024). Moreover, the design of effective prompts remains difficult and unclear. Other approaches fine-tune agents independently with individual or role-conditioned rewards. However, they require carefully curated rewards for each individual or role Slumbers et al. (2024); Liu et al. (2025a); Subramaniam et al. (2025), and, as independent learning methods, lack convergence guarantees Tan (1993).
@@
In this paper, we model LLM collaboration as a cooperative MARL problem Albrecht et al. (2024) and formalize it as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) Oliehoek and Amato (2016). In LLM collaboration, multiple trainable LLMs generate responses synchronously based on their individual prompts. The external environment evolves according to the joint responses until the dialog ends. This general model allows a wide range of problems to be modeled and solved using versions of MARL algorithms. Following the efficient practice of Group Relative Policy Optimization (GRPO) Guo et al. (2025), we propose Multi-Agent GRPO (MAGRPO) that trains LLMs in a multi-turn setting. MAGRPO leverages centralized group-relative advantages for joint optimization, while preserving decentralized execution for each agent. Our method builds off of state-of-the-art LLM approaches in GRPO and MARL approaches for centralized training and decentralized execution, such as MAPPO Yu et al. (2022). Our experiments show that MAGRPO develops various cooperation schemes, improving response efficiency with high quality.
@@
Our contributions can be summarized as follows: (i) We model the LLM collaboration as a cooperative MARL problem, where multiple LLMs cooperate to generate joint responses; (ii) We implement the MAGRPO algorithm, which optimizes agent cooperation through aligned rewards while maintaining decentralized execution to maintain efficiency; (iii) Our experiments demonstrate that fine-tuning with MAGRPO improves both response efficiency and quality in coding collaboration; (iv) We provide a detailed analysis of the limitations of existing approaches and outline open challenges in applying MARL to LLM collaboration.
2 Related Work
2.1 Test-Time Multi-Agent Interaction
@@
Recent work employs multiple agents with specialized roles interacting through diverse pipelines at test-time to enhance response quality. In multi-agent debate, agents iteratively formulate positions by reviewing other agents’ outputs, where the final decision or answer is determined by majority voting or a summarizer Du et al. (2023); Chan et al. (2023); Liang et al. (2024). Role-based approaches allocate tasks across specialized agents Wu et al. (2023a); Qian et al. (2024); Hong et al. (2024). For example, an agent may function as a verifier to assess the correctness of outputs Skreta et al. (2023); Lifshitz et al. (2025); Setlur et al. (2025), while another may act as a macro-planner to orchestrate workers’ responses. However, these multi-agent frameworks rely on prompt-level interactions among agents, often leading to ineffective communication and computational inefficiency. Moreover, the design of effective prompts and role assignment remains unclear, as prompts usually fail to reliably guide agent behavior, enforce role adherence, or support coherent coordination across tasks. These limitations motivate us to fine-tune LLMs in MAS to improve their cooperation.
2.2 Multi-Agent Fine-Tuning
@@
Recent work has explored fine-tuning LLMs to improve their performance across diverse domains, e.g., arithmetic reasoning, navigation, and hidden-role games Ma et al. (2025); Slumbers et al. (2024); Sarkar et al. (2025). These approaches typically employ individual rewards or rewards conditioned on specific roles Park et al. (2025); Liu et al. (2025a); Subramaniam et al. (2025). Such reward structures often require careful manual specification, and their underlying rationale is rarely well justified. The misaligned or conflicting incentives can hinder effective coordination. Moreover, these methods lack convergence guarantees, as each agent learns independently in a non-stationary environment where other agents are simultaneously updating their policies. In this paper, we focus on cooperative scenarios, where LLMs are jointly trained with interpretable, human-aligned rewards.
3 Cooperative MARL for LLM Collaboration
@@ 3.1 LLM Collaboration
As discussed by a number of companies NVIDIA (2024); Anthropic (2024), a team of agents could be used to generate a complex codebase. The code would be difficult, costly, and time-consuming to generate with a single agent, but a group of LLMs could do so quickly and cheaply. None of these agents is self-interested, but they are trainable in a scheme such as the one discussed below. Using a joint reward allows agents to specialize as needed to complete the task without complex prompt or reward engineering.
3.2 Problem Formalization
We formalize collaboration among LLMs as a subclass of the cooperative MARL problem, considering LLM agents and the types of problems they are solving. This problem is a form of a Dec-POMDP Oliehoek and Amato (2016), which allows cooperation through a joint reward while preserving scalable decentralized control.
@@ 3.2 Problem Formalization
In Dec-POMDP, since the states are not directly observed, each agent maintains its local observation-action history $\mathbf{h}=\{h_{1},\cdots,h_{n}\}$ to infer information about state. A solution to a Dec-POMDP is a joint policy that maximizes the expected cumulative reward, $\boldsymbol{\pi}^{*}=\{\pi_{1}^{*},\cdots,\pi_{n}^{*}\}=\arg\max_{\boldsymbol{\pi}}\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{H-1}R(s^{\text{acc}}_{t},\mathbf{a}_{t})\right]$. A joint policy is a set of local policies $\pi_{i}$, which conditions on the local observation-action history $h_{i,t}=\{o_{i,0},a_{i,0},\ldots,o_{i,t}\}$.
RL methods for Dec-POMDPs have become a popular topic (e.g., Foerster et al. (2024); Lowe et al. (2020); Foerster et al. (2018); Rashid et al. (2018); Wang et al. (2021); Yu et al. (2022); Albrecht et al. (2024)) with methods successful at scaling to large state, action and observation spaces. Many methods use Centralized Training for Decentralized Execution (CTDE), where they use some centralized information during training (e.g., a centralized value function estimate) but are still able to execute in a decentralized manner when training is complete.
3.3 Challenges in LLM Collaboration
@@
3.3.1 Representations in Natural Language
Unlike traditional cooperative MARL agents, LLM agents operate over natural language, receiving instructions and generating responses as sequences of tokens. MARL approaches could model this problem at the token or prompt/response level. At the token level, the number of actions and observations is smaller, but the problem horizon can be very long. At the prompt/response level, the actions and observations space is much larger, but the horizon is much shorter. Moreover, token-level rewards are often uninformative, as both queries and responses must form coherent and semantically meaningful structures. As adopted in prior RL methods Ouyang et al. (2022); Guo et al. (2025); Rafailov et al. (2024), we model each LLM agent’s decision-making process as a direct mapping from input instructions to complete responses to enable efficient and stable training. Nevertheless, the best modeling and solution approaches remain an open question.
3.3.2 Training Paradigm
As mentioned above, many MARL methods use centralized training for decentralized execution (CTDE). Unfortunately, standard CTDE methods use centralized value models in the form of centralized critics Foerster et al. (2024); Lowe et al. (2020); Yu et al. (2022) or mixers in value decomposition methods Rashid et al. (2018); Wang et al. (2021). Such architectures allow additional information and coordination during training but do not scale well to very large action and observation spaces (such as those in our problem). Conversely, Decentralized Training and Execution (DTE) methods Amato (2025) train a set of models, one for each agent in a decentralized manner. DTE approaches are typically more scalable but do not use additional information during training (even when it is available). It is an open question which paradigm to use to maximize performance while maintaining scalability in the LLM collaboration problem. In this paper, we balance decentralized execution with centralized training using group-based Monte Carlo estimates. Experiments show the effectiveness of our approach on short-horizon tasks.
4 MAGRPO
@@
12:   Estimate $\widehat{A}^{(g)}_{t}$, $\forall g\in\mathcal{G}$ for each branch $B$ according to Equation 1
13:   Calculate $J(\theta_{i})$, $\forall i\in\mathcal{I}$ for each branch $B$ according to Equation 2
@@
Algorithm 1 shows the procedure of MAGRPO. Given a dataset $\mathcal{D}$ containing task information (e.g., the descriptions of coding problems), $n$ LLMs are optimized, each with a policy parameterized by $\theta_{i}$ and guided by a reward model $R$. In each episode, a task is sampled from the given dataset $\mathcal{D}$, which is used to construct initial observations $\mathbf{o}_{0}=\{o_{1,0},\cdots,o_{n,0}\}$ and histories $\mathbf{h}_{0}=\{h_{1,0},\cdots,h_{n,0}\}$. Taking inspiration from the single-agent GRPO algorithm Guo et al. (2025); Liu et al. (2025b), at each turn $t$, each agent takes action by generating a group of responses $a^{\mathcal{G}}_{i,t}=\{a^{(1)}_{i,t},\cdots,a^{(G)}_{i,t}\}$ following its policy $\pi_{i}(\cdot|h_{i,t}^{\mathcal{G}})$ based on its observation-action history $h_{i,t}^{\mathcal{G}}=\{h^{(1)}_{i,t},\cdots,h^{(G)}_{i,t}\}$. The actions of individual agents are aggregated to form a group of joint actions $\mathbf{a}^{\mathcal{G}}_{t}=\{a^{\mathcal{G}}_{0,t},\cdots,a^{\mathcal{G}}_{n,t}\}$. The agents receive a group of joint rewards $r^{\mathcal{G}}_{t}$ for their responses $\mathbf{a}^{\mathcal{G}}_{t})$, which also conditions on the accessible part of the state $R(\cdot|s^{\text{acc},\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$. The joint actions triggers the transition $T(\cdot|s^{\mathcal{G}}_{t},\mathbf{a}^{\mathcal{G}}_{t})$, where agents receive new observations $o^{\mathcal{G}}_{i,t+1}=\{o^{(1)}_{i,t+1},\cdots,o^{(G)}_{i,t+1}\}$ and use them to construct histories $h^{\mathcal{G}}_{i,t+1}=\{h^{\mathcal{G}}_{i,t},a^{\mathcal{G}}_{i,t},o^{\mathcal{G}}_{i,t+1}\}$. This process continues until terminated at turn $H$.
@@
We employ stochastic gradient descent to train agents. Without explicit value models, estimating history-action values from a single rollout incurs high variance. To stabilize training, we estimate the expected return of the current state by averaging over a group of Monte Carlo samples $\{R^{(1)}_{t},\cdots,R^{(G)}_{t}\}$ for each branch $B$. As a result, we are able to generate a centralized estimate (which is common in MARL) without a large value model. For each turn $t$, the advantage of each joint action in the group is calculated as,
$\widehat{A}^{(g)}_{t}=R^{(g)}_{t}-\frac{1}{G}\sum_{g=1}^{|\mathcal{G}|}R^{(g)}_{t}$, (1)
@@
where $R^{(g)}_{t}=\sum_{\tau=t}^{H-1}r^{(g)}_{\tau}$. Inspired by GRPO Guo et al. (2025), Dr. GRPO Liu et al. (2025b), and MAPPO Yu et al. (2022), the centralized advantage values can be used to update policy $\pi_{i}$ (parameterized by $\theta_{i}$) for each agent $i$. MAGRPO does not have importance sampling and epsilon clipping since it is on-policy, and the weight of the KL divergence term is set to be 0 to encourage greater policy deviation from the base model,
@@
$J(\theta_{i})=\mathbb{E}_{\mathbf{o}_{0}\sim\mathcal{D},\mathbf{h}^{\mathcal{G}}\sim\boldsymbol{\pi}_{\boldsymbol{\theta}}}\Bigg[\frac{1}{|B|}\frac{1}{|\mathcal{G}|}\sum_{h_{i}^{\mathcal{G}}\in B}\sum_{g\in\mathcal{G}}\widehat{A}^{(g)}_{t}\log\pi_{\theta_{i}}(a_{i,t}^{(g)}|h_{i,t}^{\mathcal{G}})\Bigg],$ (2)
where the gradient is averaged across all the branches (states) and over the group of generations at $t$.
5 Experiments
(a) MAGRPO-ST on HE
(b) MAGRPO-MT on HE
(c) MAGRPO-ST on CHE
(d) MAGRPO-MT on CHE